### v1

# Float Self-Tagging

Abstract.
Dynamic and polymorphic languages must attach information, such as types, to run time objects, and therefore adapt the memory layout of values to include space for this information. This is especially problematic in the case of IEEE754 double-precision floating-point numbers, which require exactly 64 bits, leaving no space for type information. The two main encodings in-use to this day, tagged pointers and NaN-tagging, either allocate floats on the heap or unbox them at the cost of an overhead when handling all other objects.
This paper presents self-tagging, a new approach to object tagging that can attach type information to 64-bit objects while retaining the ability to use all of their 64 bits for data. At its core, self-tagging exploits the fact that some bit sequences appear with very high probability. Superimposing tags with these frequent sequences allows encoding both 64-bit data and type within a single machine word. Implementations of self-tagging demonstrate that it unboxes all floats in practice, accelerating the execution time of float-intensive benchmarks in Scheme by 2.3$\times$, and in JavaScript by 2.7$\times$ without impacting the performance of other benchmarks, which makes it a good alternative to both tagged pointers and NaN-tagging.
Dynamic Languages, Polymorphic Languages, Floating-Point, NaN-Tagging, Compiler Optimization, JavaScript, Scheme
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Software and its engineering Source code generationccs: Software and its engineering Polymorphism
1. Introduction
In dynamic and other polymorphic languages, efficiently handling floating-point numbers (floats) remains an elusive problem in search of an optimal solution. These languages require attaching type information to values at run time, including numeric types like integers and floats. This requirement conflicts with the IEEE754 standard (iee, 2019) for encoding floats, which uses a full 64-bit word to represent a double-precision float, leaving no space for additional type information. Various representations have been developed to mitigate this issue, each with its own trade-offs.
This paper presents a novel technique for encoding floats that avoids most float allocations without adversely affecting the cost of checking types and accessing objects, which are frequent operations of dynamic languages. This technique does not require any programs static analysis and could be integrated in most runtime systems of languages that need run time types. It is developed in the context of 64-bit runtime systems but also has applications on 32-bit systems. The rest of this section presents the main encodings in-use as well as their trade-offs.
1.1. Tagged Objects
The classical and popular solution to preserve type information is to attach a tag to all objects (Gudeman, 1993; Leijen, 2022; Jones et al., 2012a; Appel, 1998). A machine word is used to encode an object reference (either a value or a pointer to a heap-allocated value) and a short sequence of bits is reserved in the word to store type information. For the rest of the paper the term object is used interchangeably with object reference. Tagging allows low-cost type checks and object access at the cost of losing a few bits for data. Aligning all heap-allocated values to 64-bit machine words conveniently frees the low bits of pointers to store a 3-bit tag. On many architectures, accessing the fields of the object can be done at no additional cost by using an offset in the memory dereference instruction to take into account the tag for that type of object. Objects can thus be encoded using one of the three following representations.
  • Tagged values store type information with a tag, and the object’s value in the remaining bits. This representation does not require heap allocation.
  • Tagged pointers also store type information with a tag, but values are stored in the heap. The objects are represented by pointers with a tag in their low bits.
  • Generic pointers represent all remaining object types as pointers with a generic tag. Since the generic tag is associated with multiple types, type information is stored in memory, typically in a header.
Hence, a tag indicates a type, but also a memory representation. A tagged value representation is the most efficient as it needs no heap allocation and no memory read. A tagged pointer still allows efficient type checks, but requires heap allocation. This introduces the overhead of dereferencing the pointer, and adds a strain on the garbage collector. A generic pointer is the least efficient representation as it involves storing type information in memory, which increases space usage and the cost of type checks and memory management. Consequently, specific tags are conscientiously assigned to the most frequent types observed in programs, such as small integers or floats. Figure 1 shows a typical memory layout of each representation.
Figure 1. The three representations in a tagged object system (here shown on a little-endian machine)
As an example, consider an implementation using 3-bit tags. Due to the frequent use of small integers, a tagged value representation is appropriate for small integers and the choice of the tag 000 allows addition/subtraction directly on the tagged values. Since 3 bits are occupied by the tag, small integers are limited to 61 bits. This is generally considered an acceptable trade-off since a two’s complement representation of signed integers can accommodate any number of bits, and 61 bits still offers a wide range of values. When larger integers are required, generic pointers to heap-allocated big integers can be used (Leijen, 2022). Consequently this provides a hybrid representation where the most common integers appear as efficient tagged values and less-frequent big integers are represented with more costly generic pointers.
Unfortunately, when it comes to the representation of floats, none of the aforementioned options is well-suited. Double-precision floats cannot be represented as tagged values since the IEEE754 standard enforces the use of 1 sign bit, 11 exponent bits and 52 mantissa bits (see Figure 2(a)), totaling 64 bits and leaving no space for a tag (iee, 2019). Parting from this standard is impracticable due to the set of instructions offered by X86 (Corporation, 2024) and ARM (Limited, 2009) architectures being specific to 32 or 64-bit floats.
Yet, tagged pointers introduce two major inefficiencies for floats. First, all uses of the value require fetching it from memory. Second, the heap allocation of all floats increases memory usage and the cost of computing float results. This is especially disconcerting considering that allocated floats are often short-lived intermediate results that add a strain on garbage collectors (Serrano and Feeley, 1996; Jones et al., 2012b). This has costly implications for programs performing extensive numerical computations or languages such as JavaScript where floats are the only available primitive number type (International, 2015).
Still, tagged objects are straightforward to implement and are thus found in numerous compilers such as V8 (Sheludko and Solanes, 2020), QuickJS (Bellard and Gordon, 2024), and Hopc (Serrano, 2021) (JavaScript), the Lua interpreter (Roberto Ierusalimschy, 2024) (Lua), CRuby (Shaughnessy, 2013) (Ruby), SBCL (Newman et al., 2024) (Lisp), and Bigloo (Manuel Serrano, 2024) (Scheme). Recently, CPython also moved toward tagged pointers in the process of removing its global interpreter lock (Gross, 2023).
1.2. NaN-tagging
NaN-tagging circumvents the drawbacks of tagged pointer floats by reclaiming unused bits in the encodings of floating-point NaN values to store data (Gudeman, 1993; Leijen, 2022). As per the IEEE754 standard, a NaN value is represented by setting all 11 bits of the exponent to 1 and a non-zero mantissa (zero is used to encode Infinity). The mantissa’s highest bit distinguishes between a quiet and signaling NaN, which determines whether the NaN should signal an exception or fall through operations.
A subset of NaN’s can thus be reserved to encode non-float objects. A convenient subset is that of negative, quiet NaN’s, illustrated in Figure 2(b), which correspond to the interval from 0xfff8:$0$:$0$:$0$ to 0xffff:ffff:ffff:ffff. This partitions floating-point numbers in two intervals. Sequences of bits above 0xfff8:$0$:$0$:$0$ are reserved for non-float objects and those lesser than or equal to 0xfff8:$0$:$0$:$0$ are reserved for floats with 0xfff8:$0$:$0$:$0$ itself representing a negative, quiet NaN.
(a) IEEE754 floating-point representation of $\pi$.
(b) NaN-tagging uses the 51 free bits on negative, quiet NaN values to store data.
Figure 2. Floating-point representations of numbers and NaN.
On current hardware, memory addresses typically fit in 48 bits. The 51 bits uncovered by NaN-tagging are thus sufficient for storing tagged pointers with 3-bit tags, and 32-bit small integers. It offers the advantage of unboxed floats, but impacts the performance of all other, non-float objects due to higher-cost machine instructions necessary to check the type and dereference NaN-tagged pointers.
This reliance on hardware specific details also interferes with other optimizations and portability. For instance, NaN-tagging prevents (or largely complicates) optimizations relying on stack allocations. The stack uses high memory addresses that do not fit in 48 bits unless encoded relative to the location of the stack segment. Another caveat of NaN-tagging is the impossibility to apply it on 32-bit architectures as 32-bit NaN’s would severely cripple memory management by leaving only 22-bit pointers, which only allows addressing about 4 MB.
NaN-tagging is used in a few language implementations, including SpiderMonkey (Foundation, 2024) (JavaScript), tinylisp (Engelen, 2024) (Lisp), LuaJIT (Pall, 2024) (Lua), and Zag (Mason, 2022) (Smalltalk).
1.3. Pointer-biased NaN-tagging
NuN-tagging is a variant of NaN-tagging that alleviates the cost of dereferencing NaN-tagged pointers (Leijen, 2022). It relies on the (currently valid) fact that common hardware never returns negative, quiet NaN’s greater than fffd:ffff:ffff:ffff.
Figure 3. Encoding of Nun-tagged values, after applying the 0x0001:$0$:$0$:$0$ offset. The low bit sequences are used for non-float objects that can be represented as tagged pointers or value.
Instead of taking advantage of the unused range of bit sequences from 0xfffe:$0$:$0$:$0$ to 0xffff:ffff:ffff:ffff, which would result in an encoding similar to standard NaN-tagging, NuN-tagging offsets all floats by adding 0x0001:$0$:$0$:$0$. This frees up the lower bit sequences for 48-bit tagged objects as shown in Figure 3. It however reintroduces a cost for encoding and decoding floats to operate on their value.
NuN-tagging is used by JavaScriptCore (Inc., 2024) (JavaScript).
1.4. Contribution
The contribution of this paper is a new approach to object tagging, named self-tagging, that allows attaching type information to some 64-bit objects while retaining the ability to use all of their 64 bits for storing data. This approach is applied to implement double-precision floats as tagged values instead of tagged pointers. Contrary to NaN-tagging, the newly obtained representation does not impact the performance of encoding and decoding other tagged objects and it is compatible with stack addresses. It is also applicable in the context of single-precision floats on 32-bit architectures, which NaN-tagging does not support. In practice, self-tagging is shown to prevent the heap-allocation of all floats, thus providing the advantages of both tagging and NaN-tagging without their main drawbacks.
1.5. Paper Structure
This paper is structured as follows. Section 2 describes the general idea of self-tagging with a focus on its application to floats. Section 3 provides a first implementation of self-tagging in the Bigloo Scheme compiler. Section 4 shows how to refine the technique to unbox all floats in the Bigloo compiler and in the Hopc ahead-of-time JavaScript compiler. Section 5 presents technical details of self-tagging implementation. It shows portable implementations that could be used by compilers that generate either C or assembly code. Section 6 discusses the adaptation of self-tagging to 32-bit architectures. Finally, section 7 presents related work.
2. Self-Tagging
This section describes self-tagging, a new tagging technique that exploits the fact that some values naturally contain the appropriate tag corresponding to their type, at the correct location, in their bit arrangement. While this may appear coincidental, it occurs frequently enough to allow the transformation of a significant proportion of heap allocated objects (either represented as tagged pointers or generic pointers) to tagged values. In this paper, unless stated otherwise, double-precision floats and a 64-bit architecture are implied. However, self-tagging pertains to 32-bit architectures as well, which is discussed in Section 6.
Consider the common case of tagged objects with 3-bit tags and, for the sake of the example, assume that the tag 001 is chosen for floats. This setup was previously illustrated in Figure 1 where a tagged pointer refers to the heap-allocated float 3.02. The IEEE754 encoding of 3.02 happens to end with 001, the tag for floats in this example. This purely coincidental fact is bound to happen with $1/8$ of floating-point representations, since eight distinct tags exist on 3 bits.
Self-tagging exploits such occurrences where the tag bits of a pointer appear in its value. Such objects can be unboxed, making them tagged values instead of heap allocated value. However, since only $1/8$ of all floats can be unboxed in such a way, a second tag must be reserved for the remaining floats, which still need to be represented as heap allocated values with either tagged or generic pointers.
Self-tagging introduces the notion of tags that indicate which values are unboxed. This contrasts with standard object tagging that invites assigning tags with efficient handling of frequent types in mind, regardless of values. With self-tagging, the initial choice of 001 for floats ought to be reevaluated. A better choice could be 000, which captures, among other values, all integer valued floats below $12^{50}+$ 1 (the first integer valued float whose IEEE754 representation does not end in 000). The tag 001 can be kept for tagged pointer floats.
Figure 4. Representation of a (unboxed) float as a tagged value with tag 000. All floats whose IEEE754 representation ends with 000 are unboxed. Other floats are encoded as tagged pointers as usual.
Figure 4 shows the encoding of floats whose IEEE754 representation ends with the better suited tag 000 for unboxed floats. In those cases, the last three bits of the mantissa and the tag are identical. This allows superimposing them to represent both a 64-bit value and a 3-bit tag within a single machine word. Any object tagged with 000 is known to be a tagged value float. Obtaining its value requires no further transformation.
This specific example of self-tagging allows unboxing $1/8$ of floats. Yet, this proportion can be increased by assigning extra tags to unboxed floats. In general, if $T$ tags are available and $f$ tags are reserved for self-tagging of floats, then $f/T$ of all floats can be unboxed. A higher number of reserved tags $f$ can potentially lead to less memory allocation, and thus less strain on the garbage collector. However, the short supply of tags suggests that $f$ should be limited and only worthwhile tags should be selected. Tag selection is discussed in Section 4.
3. A First Experiment
This section presents the first experiment of float self-tagging with the Bigloo Scheme compiler (Manuel Serrano, 2024). Bigloo uses tagged objects by default and can be configured to use NaN-tagging (discussed in Section 4.2), making it ideal for comparing these techniques and self-tagging.
3.1. Overview of the Bigloo Compiler
The Bigloo compiler builds a typed abstract syntax tree from the compiled Scheme program. Several analysis and optimization refine the initial optional annotations contained in the source program, the main ones being occurrence-typing (Tobin-Hochstadt and Felleisen, 2010) and storage use analysis (Serrano and Feeley, 1996). During compilation, primitive values have a dual representation: one for polymorphic contexts and one for specific contexts. For instance, when it can establish that a function is always invoked with floats, the compiler assigns the C double type to that function and, when needed, it introduces cast operations from and to the generic representation of floats and double.
The result of this compilation technique is that most temporary and local functions are precisely typed and avoid polymorphic representations, which contributes to removing otherwise necessary heap allocations and casts for floats. The compiler also tracks cases where only floats are stored into vectors, which are then transformed into vectors of C doubles. The experiment conducted in this section evaluates the impact of removing the memory allocations of floats that the compiler has not been able to remove statically. For a compiler that does not implement such static analyses, the impact might be even more significant than measured with Bigloo.
3.2. Experimental Setup
Two versions of the Bigloo Scheme compiler are compared. The first is the original Bigloo version, which implements standard tagged objects where floats are heap-allocated using a tagged pointer. The second implementation only differs by its use of self-tagged floats.
In the self-tagging variant, the two tags 000 and 100 are reserved for self-tagged floats, thus increasing the proportion of self-tagged floats to $1/4$. This means that all floats $n$ that are integers such that $|n|\leq 2^{51}$ as well as those values scaled by a power of two, will be unboxed floats. The tag 001 is reserved for the remaining floats, which are represented by tagged pointers.
Both implementations are tested using a subset of the R7RS benchmarks, which is the standard benchmark suite for Scheme (r7r, 2024). The experiment includes all macro-benchmarks (more than 500 lines of code) and benchmarks whose calculations mainly involve floats (tagged as such in the R7RS benchmarks suite). For the purpose of analyzing results, it is useful to distinguish between float-intensive benchmarks (fibfp, fft, mbrot, nucleic, pnpoly, ray, simplex, sum1, and sumfp) and other benchmarks that use few of no floats.
The following results were obtained on a machine with an Intel Xeon W-2245 CPU 3.90GHz and 32 GB of RAM, under Debian 13 with Linux Kernel 6.10.6.
3.3. Memory Allocation
Memory usage is measured by counting the number of heap-allocated bytes across the lifetime of each benchmark. The original and self-tagging Bigloo versions are then compared to assess the efficiency of self-tagging at unboxing floats and its impact on benchmarks that make limited use of floats. Results are shown in Figure 5.
Figure 5. Heap allocations comparison between the original Bigloo version and the self-tagging version. The y-axis shows heap allocations relative to the original version. 1.0 indicates no change, while 0.0 indicates that all allocations were removed.
On all benchmarks extensively using floats, memory allocation decreases. On mbrot, nucleic, simplex and sum1, allocations decreases by about 25%, which is consistent with the proportion of self-tagged floats ($1/4$). The fft benchmark has only 17% fewer allocations due to predominantly using the float 0.0, which the original Bigloo version preallocates. For fibfp, pnpoly, ray and sumfp there is very little allocation of floats when using self-tagging because those programs mostly use “simple” floats that are integer valued floats $\leq 2^{51}$ scaled by a small power of two, such as 0.75, 2.25, or 42.0.
On all benchmarks that use few floats, memory usage increases by less than 1%. This tenous increase stems from the use of generic pointers for types that were assigned specific tags in the original version.
3.4. Execution Time
Figure 6 shows the execution time comparison between both implementations. Execution time refers to elapsed real time measured with perf-stat. Each benchmark is parameterized for a run to last for at least 10 seconds, and is repeated 5 times. All timing results are the average execution time of these runs.
Figure 6. Execution time comparison between the original Bigloo version and the self-tagging version. The y-axis shows execution time relative to the original version. 1.0 indicates no change. Lower means faster execution than the original version.
On a geometric average, float-intensive benchmarks are about 25% faster while non-float benchmarks are unaffected. However, the relation between lower memory usage and faster execution is noisy. The execution of mbrot and nucleic is significantly slower despite allocating about $1/4$ less floats with self-tagging.
This unpredictability is explainable, and fixable. Figure 7 shows that, despite reducing float allocations in mbrot and nucleic, self-tagging also causes about 6 times more missed branch predictions. This stems from the fact that tags occupy the low bits of the mantissa. In that position, the smallest variations cause a float’s representation to switch between tagged value and tagged pointer in a hard to predict way. This is beyond the capabilites of the branch predictor and execution suffers costly misprediction penalties (Bryant and O’Hallaron, 2016).
Figure 7. Missed branch predictions comparison between the original Bigloo version and the self-tagging version. The log-scaled y-axis shows relative number of missed branch predictions compared the original version. 1.0 indicates no change, lower means fewer missed predictions and higher means more missed predictions.
4. Improved Predictability and Coverage
This section addresses the issue uncovered by the previous experiment and presents an improved, performant implementation of float self-tagging.
The previous experiment shows that while self-tagging decreases the number of heap-allocated floats it does not consistently improve performance due to failed branch predictions. To fix this issue, the tag system can be modified to bias it in favor of common floats by superimposing the 3 tag bits with the most significant bits of the exponent, as shown in Figure 8. Hence, all floats whose exponent falls into a given range, which is determined by the tag, are unboxed. Since tags must be placed on low bits, floats are encoded to tagged values by a 60-bit rotation to the right (or 4-bit to the left). Decoding is achieved with the inverse rotation.
Figure 8. A float self-tagging representation where the tag corresponds to the high bits of the exponent. The top bit sequence is a float where the tag is superimposed with the exponent’s three most significant bits. The bottom sequence is the tagged value representation of the float where a 4-bit left rotation is applied to place the tag on low bits.
Given a tag, the floating-point intervals it covers can be computed as follows. Assuming a 0 sign bit, the lowest and highest bounds are reached when all eight exponent low bits and all mantissa bits are set to 0 and 1 respectively. There are two exceptions when the tag is either 000 or 111. The former represents subnormal numbers when all exponent bits are set to 0, in which case the lowest bound is 0.0. In the case of the tag 111, when all exponent bits are set to 1, the floating-point representation is that of Infinity or NaN. Since the sign bit is not part of the tag, each tag also captures a corresponding negative interval.
Exponent Value
Tag range range fft fibfp mbrot nucleic pnpoly ray simplex sumfp sum1
0.0 94% 13% 0% 0% 16% 13% 11% 0% 0%
000 0..255 $2^{-1074}$..$2^{-767}$
001 256..511 $2^{-767}$..$2^{-511}$
010 512..767 $2^{-511}$..$2^{-255}$
011 768..1023 $2^{-255}$..2222 6% 33% 96% 70% 60% 18% 62% 0% 0%
100 1024..1279 $2$..$2^{257}$ 0% 54% 4% 30% 24% 69% 27% 100% 100%
101 1280..1535 $2^{257}$..$2^{513}$
110 1536..1791 $2^{513}$..$2^{769}$
$2^{769}$..$2^{1024}$
111 1792..2047 Infinity/NaN - - - - - - - - -
(a) The positive value ranges captured by each tag and the proportion of all floats generated by the R7RS float-intensive benchmarks. The lowest and highest exponents are exponents before applying the bias ($-1023$ for 64-bits floats). Values in the interval formed by the lower bound (included) and upper bound (excluded) can be unboxed by self-tagging for the given tag. Intervals of interest are highlighted. Empty entries indicate no float in this range was generated, whereas 0% means that very few floats in this range were generated (less than 0.5%).
(b) Logarithmic scale number line of float values, excluding NaN values and infinities. Intervals of interest are highlighted with their corresponding color from Figure 9(a).
Figure 9. Floating point interval captured by each tag.
Figure 9(a) shows the positive interval captured by each tag. Three tags stand out as capturing worthwhile intervals. The tags 011 and 100 capture the range of floats from $2^{-255}$ to $2^{257}$, or about $1.7\times 10^{-77}$ to $2.3\times 10^{77}$. The tag 000 captures the values $\pm$0.0 and, incidentally, the floats between $-2^{-767}$ and $2^{-767}$. Figure 9(b) shows the intervals captured by these tags on a number line.
Contrarily to the implementation of self-tagging in Section 3, these tags capture intervals of contiguous values. This contiguity within intervals prevents small variations from causing switches from a tagged value to a tagged pointer representation.
In fact, although the tagged floats cover only $3/8$ of all floats, it is reasonable to expect that many programs will use floats in these intervals with very high probability, which benefits branch prediction. The right part of Figure 9(a) confirms this expectation. It presents the distribution of floats that are generated in each float-intensive R7RS benchmark. Values are aggregated by tag, with the exception of $\pm$0.0 and Infinity/NaN, which are counted separately. The tags 000, 011 and 100 cover all floats in these benchmarks. Moreover, among the values captured by 000, only $\pm$0.0 are used.
This permits dropping the tag used for the tagged pointer representation of the remaining floats due to their expected rarity. They are instead represented with generic pointers with a header type. Whether expending the tag 000 to capture 0.0 and -0.0 is a desirable trade-off is discussed in the next section.
4.1. A Second, Improved Experiment
The experimental setup from Section 3 is replicated to test float self-tagging using high bits of the exponent. Two variants are compared to the original version of Bigloo that implements standard tagged objects.
  • self-tagging (3 tags) uses the tags 000, 011 and 100, thus unboxing floats in the intervals from Figure 9(b). Remaining floats are represented by generic pointers with a header type.
  • self-tagging (2 tags) tests the impact of expending the tag 000 for self-tagging. It only uses the tags 011 and 100 for float self-tagging and reclaims the tag 001 to represent all remaining floats as tagged pointers. The floats $\pm$0.0 are preallocated.
Figure 10 shows the memory allocations of the improved variants. Both variants allocate fewer floats on all float-intensive benchmarks. In fact, both versions have been verified to allocate no float at all (see Figure 9(a)) with the only exceptions being the preallocated $\pm$0.0 by the second version. The remaining heap usage is for non-float objects. Similarly to the previous experiment (Figure 5), memory usage of other benchmarks remains mostly unaffected.
Figure 10. Heap allocations comparison between both versions with self-tagging on exponent bits. The 3 tags version uses tags 000, 011 and 100. The 2 tags version uses only 011 and 100, but preallocates 0.0 and -0.0. The y-axis shows heap allocations relative to the original Bigloo version, which uses tagged pointers for all floats. 1.0 indicates no change, while 0.0 indicates that all allocations were removed.
Figure 11 shows faster execution times for both variants of self-tagging. All float-intensive benchmarks execute faster than the original Bigloo implementation, with the exception of fft on the version that does not use the tag 000. In particular, the mbrot and nucleic benchmarks are both faster, confirming that self-tagging on exponent bits eliminates the branch misprediction penalty incurred by self-tagging on mantissa low bits (see Section 3). Despite small individual variations, other benchmarks remain generally unaffected.
The version using three tags for self-tagging, including the tag 000, is 2.3$\times$ faster on average than the original Bigloo version. This is significantly higher than the 1.7$\times$ acceleration without using the tag 000. In particular, the benchmark fft, which predominantly uses the values $\pm$0.0, is slower in the second version despite zeros being preallocated. Hence, it is better performance-wise to include the tag 000.
Figure 11. Execution time comparison between both versions with self-tagging on exponent bits. The 3 tags version uses tags 000, 011 and 100. The 2 tags version uses only 011 and 100, but preallocates 0.0 and -0.0. The y-axis shows execution time relative to the original version. 1.0 indicates no change, lower means faster execution than the original Bigloo version, which uses tagged pointers.
4.2. Application to JavaScript
This section experiments with the application of float self-tagging in JavaScript, which is a prime candidate for self-tagging because it has no integer type. Its specification defines numbers as double-precision floats (International, 2015). To isolate the impact of the representation of float, three versions of the Hopc JavaScript compiler are tested, where each version differs only by its representation of floats, which is either tagged pointers, NaN-tagging or self-tagging.
Hopc is an ahead-of-time JavaScript compiler that uses Scheme as its compilation backend and relies on the underlying Scheme compiler for optimizations (Serrano, 2021). This permits reusing Bigloo’s runtime to test float self-tagging in the context of JavaScript. Bigloo can also be configured to use NaN-tagging. This configuration is scarcely used for Scheme because of its negative impact on other types, but is applicable to compilation of JavaScript.
Figure 12 compares Bigloo’s NaN-tagging performance to that of float self-tagging with three tags from Section 4.1. On float-intensive benchmarks, self-tagging generally fares slightly better than NaN-tagging. Moreover, on non-float benchmarks the float-bias of NaN-tagging makes it slower than the original implementation on average, while the self-tagging version remains generally unaffected.
Figure 12. Execution time comparison between Bigloo with self-tagging on exponent bits and Bigloo with NaN-tagging. The y-axis shows execution time relative to the original Bigloo version, which uses tagged pointers. 1.0 indicates no change, lower means faster execution than the original version.
Figure 13 presents the results of porting Bigloo’s NaN-tagging and self-tagging implementations to JavaScript. Performance are measured using the Hop benchmark suite, which gathers benchmarks from the popular JetStream, Octane and Sunspider JavaScript benchmark suites (Serrano, 2024). Once again, it is useful to categorize benchmarks between float-intensive benchmarks (almabench, n-body, navier-stokes, and raytrace), and benchmarks that use few or no floats.
The Hopc NaN-tagging and self-tagging implementations have comparable performance on float-intensive benchmarks. Execution time with NaN-tagging is faster on almabench and n-body, but slower on raytrace, and equivalent on navier-stokes. However, on other benchmarks the self-tagging implementation generally fares better due to its lower overhead when handling non-float objects.
Figure 13. Execution time comparison between Hopc with self-tagging on exponent bits and Hopc with NaN-tagging. The y-axis shows execution time relative to the original Hopc version, which uses tagged pointers. 1.0 indicates no change, lower means faster execution than the original version.
5. Implementation
This section discusses the implementation of the float-related operations of the dynamic language. The type f64 will refer to IEEE754 64-bit doubles and the type i64 will refer to 64-bit integers. Moreover, the type object will refer to the values of the language (the sum type of floats, integers, booleans, and other objects), which is an i64. To better understand the low-level implications of the implementation, assembly code required on x86-64 is shown to illustrate the details of the implementation on that architecture.
5.1. Tag Testing
Testing the low 3-bit tag of an i64 value is a basic need for float-related operations and dynamic type checks. Checking that the 3 low bits are equal to a specific tag can be achieved by masking those bits followed by a comparison. Assuming the i64 value is in the x86 64-bit register rax, then the following sequence will branch to label matching_tag when the 3 low bits are equal to tag:
and al,#3
cmp al,#tag
jz matching_tag
Note that by using al only the lower 8 bits of rax are accessed, which leads to smaller constants and instruction encodings. Unfortunately, this sequence modifies rax so if the value is needed after the test, as often will be the case, then an additional instruction and register are required to keep a copy. In the special case of tag=0, the two first instructions can be replaced with “test al,#7”, which also eliminates the need for a copy because it does not modify rax.
Because a tagged float will have one of a set of tags, it is useful to check this more efficiently than a sequence of comparisons. This can be done with the bit-test instruction “bt n,i” that reads the bit at index i of n and puts it in the carry flag. The bt instruction comes in 16, 32, and 64 bit variants, but not 8 bit. As an example, testing that the least 3 low bits of rax are either 000 (0), 011 (3), or 100 (4), can be done with:
mov bx,#0x1919 # set of tags
bt bx,ax # test bit at index ax of bx
jc matching_tag
The bit index in ax is obtained mod 16, so the byte 0b00011001=0x19 (all 0’s except at bit indices 0, 3, and 4) is repeated twice in register bx. Using the bt instruction has the advantage of neither modifying the rax register nor the register holding the set of tags, so the cost of the mov instruction can be amortized over multiple checks of that set of tags (a possibly near zero cost if that register is globally reserved).
On architectures without a bit test instruction, a dynamic count shift of the tag set register could be used to achieve the bit indexing. If this is done by modifying the tag set register the initialization can’t be shared by multiple tag tests. On most 3-address RISC architectures the shift can be done non-destructively. On ARM A64, here is a short instruction sequence that tests the 3 low bits of 64-bit register x1 for a tag of 0, 3, or 4:
lslv w3, w2, w1
cmp w3, 0
b.mi matching_tag
It assumes that the 32-bit register w2 has been preloaded with 0x98989898, the bit reversed tag set that aligns the bit for tag 0 with the sign bit.
The cost of the check does not depend on the number of tags tested. This can be advantageous if a tagged pointer is used for boxed floats, say with tag b. In that case, a check for a float (either tagged or boxed) can be done by adding b to the tag set. Once it is known that an object is a float, it is easy to check for the single tag b to discriminate between tagged and boxed representations.
[…]
To convert a f64 to an object, it must first be converted to an i64, and then the high bits of the exponent must be aligned with the 3 low bits using a rotate instruction. If those 3 low bits are one of 000, 011, or 100 then the conversion is done. Otherwise, an out of line routine can be called to allocate a boxed float and return an appropriately tagged pointer. If the f64 value is in the 64-bit float register xmm0, then the following x86 code will set register rax to the corresponding object:
mov rax,xmm0 # rax $\leftarrow$ xmm0
rol rax,#4 # rotate left 4 places
bt bx,ax # assumes bx initialized to 0x1919 elsewhere
jc done
call allocate_boxed_float
done:
Assuming the initialization of the tag mask register is amortized over multiple float-related operations, the cost for the hot path is low: 3 register-to-register operations and an easily predictable conditional jump.
[…]
To convert an object (that is known to be a float) to a f64, the tag must be tested to see if it is a tagged float. If it is, rotating the bits 4 places to the right will restore the bits of the f64 result. Otherwise, a memory read gets the f64 result out of the boxed float. If the object value is in the 64-bit register rax, then the following x86 code will set register xmm0 to the corresponding f64:
bt x,ax # assumes bx initialized to 0x1919 elsewhere
jc rotate
mov xmm0,[rax+offset] # xmm0 $\leftarrow$ value field of boxed float
jmp done
rotate:
ror rax,#4 # rotate right 4 places
mov xmm0,rax # xmm0 $\leftarrow$ rax
done:
Here too the cost is low for the hot path: 3 register-to-register operations and an easily predictable conditional jump.
5.4. C Implementation
This section explains how the above operations can be implemented in C, as this is a common implementation language, in particular it is used by the Bigloo Scheme to C compiler used in the experiments of Sections 3 and 4.
Checking for a set of tags can be implemented with the following pure C function, which is easily inlined by a C compiler:
#define TAG_SET ((1<<4)|(1<<3)|(1<<0)) /* 0x19 for tags 100, 011, and 000 */
inline bool has_tag_4_or_3_or_0(int64_t n) {
return (((uint32_t)1 << n) & (~(uint32_t)0/0xff * TAG_SET)) != 0;
The expression (uint32_t)1<<n implicitly computes a shift count mod 32 and the bitwise-and therefore tests the bit at index n mod 32 of its second operand. The code uses the compile-time constant ~(uint32_t)0/0xff * TAG_SET that repeats the 8 bit mask TAG_SET 4 times to fill a 32 bit word, giving 0x19191919. This achieves the equivalent of testing the bit at index n mod 8 of TAG_SET.
Using this pure C definition with GCC version 13.2.0 and clang version 18.1.3 on x86-64 results in both cases in the non-destructive approach based on the bit test instruction. On older versions of those compilers that would implement this with a destructive shift instruction, the bit test approach can be achieved using the following definition that uses an asm statement:
inline bool has_tag_4_or_3_or_0(int64_t n) {
bool carry;
__asm__("bt%%ax,%2;" /* get one bit of mask into carry (ax is index) */
: "=@ccc"(carry)
: "a"((uint16_t)n),
"r"(~(uint16_t)0/0xff * TAG_SET)); /* 0x1919 */
return carry;
Boxing and unboxing operations need to convert between a float value and its bit representation. This can be achieved portably with a union type whose fields are of type f64 and i64. Moreover, boxing and unboxing need to rotate the bit representation of the float. Although C does not provide an operator for bit rotation, both GCC and clang recognize an equivalent pair of shifts and generate a single machine rotate instruction. The implementation in C is:
union di { double d; int64_t i; };
#define ROTR(n,s) ((int64_t)(((uint64_t)n >> s) | ((uint64_t)n << (64 - s))))
inline int64_t f64_to_object(double f) {
int64_t result = ROTR(((union di)f).i, 60);
if (has_tag_4_or_3_or_0(result)) return result;
return allocate_boxed_float(f);
inline double object_to_f64(int64_t o) {
int64_t result = ROTR(o, 4);
if (has_tag_4_or_3_or_0(result)) return ((union di)result).d;
return value_of_boxed_float(o);
6. Self-tagging on 32-bit architectures
Although previous sections only considered double-precision floats on 64-bit architectures, it is straightforward to adapt float self-tagging to single-precision floats on a 32-bit architecture. The core idea, which is to superimpose a tag to a sequence of bits likely to appear in practice, can be adapted to any machine word size.
On 32-bit implementations, it is frequent to have four 2-bit tags and heap-allocated objects aligned to 32-bit words. One option is to select two tags for float self-tagging, one tag for small integers, and assign the remaining tag to generic pointers. A similar interval analysis as in Section 4 leads to the choice of 01 and 10 which capture the interval from $2^{-63}$ (included) to $2^{65}$ (excluded), roughly $1.1\times 10^{-19}$ to $3.7\times 10^{19}$, and the equivalent negative interval illustrated in Figure 14.
Figure 14. Logarithmic scale number line of single-precision float values unboxed by self-tagging.
This tag selection covers half of all floats, but does not capture the values $\pm$0.0. However, the experiment from Section 4.1 demonstrated that, by preallocating $\pm$0.0, it can improve performance.
As a final remark, self-tagging can be applied in an implementation using two 1-bit tags with a single tag that captures half of all floats (provided the only free tag is not assigned to small integers). In this context, one may consider superimposing the tag with the float sign bit to capture all positive floats.
7. Related Work
Unboxing floats is a longstanding problem in dynamic language implementation. In the last decades, little progress has been made improving the encoding of floats, with most implementations either using boxed floats or suffering from the overhead of NaN-tagging on pointers. Thus, more general strategies that use data flow analysis (Henglein, 1992; Serrano and Feeley, 1996; Petersen and Glew, 2012; Lindahl and Sagonas, 2003) and partial evaluation (Würthinger et al., 2017) have been developed to find locations where the type of a value is known across its lifecycle. This allows the generation of specialized code that safely handles fully unboxed, untagged values. Such strategies are not specific to unboxing floats. Rather, they tackle the more general problem of inferring types in dynamic languages, which allows dropping type information in either tagged pointers or NaN-boxed pointers. While more limited in scope, self-tagging solves the problem of boxed floats with a new encoding that involves no program analysis.
Self-tagging is straightforward to add to compilers that already use object tagging. Therefore, implementations that represent floats as tagged pointers could benefit from it with minimal implementation effort. Such popular implementations include CPython (Gross, 2023), CRuby (Shaughnessy, 2013) and Google’s V8 (Sheludko and Solanes, 2020). Compilers that use NaN-tagging, such as Mozilla’s SpiderMonkey (Foundation, 2024) and Apples’s JSC (Inc., 2024) can also benefit from the lower impact that self-tagging has on the performance of non-float types compared to NaN-tagging.
Due to the shortcomings of tagged pointers and NaN-tagging, additional strategies are used to avoid boxing. A straightforward approach is to provide homogeneous data structures such as TypedArray in JavaScript (International, 2015) and numpy’s float64 arrays in Python (Harris et al., 2020). Type homogeneity then allows unboxing values within the collection. More generally, compilers can detect data homogeneity at run time and use context-dependent storage strategies to store data in a way that prevents pointer chasing (Bolz et al., 2013). Since self-tagging unboxes all floats in practice, it effectively prevents such pointer chasing in the case of floats without static of run time analysis.
In this paper, self-tagging was discussed primarily in the context of dynamic languages. Yet, polymorphic languages such as OCaml and Haskell also attach information to run time values and spend considerable effort to find efficient encodings for abstract data types (Baudon et al., 2023; Kennedy and Vytiniotis, 2012). Self-tagging could be applied in these languages, either to floats or to the encoding of abstract data types.
8. Conclusion
This paper has presented a new approach to object tagging that allows attaching type information to 64-bit values at low cost. The core idea is to superimpose tags to bits that appear with high probability in practice. This approach is especially useful to represent IEEE754 floating-point numbers, whose encoding requires all bits of a 64-bit machine word.
In practice, self-tagging unboxes all floats, reducing the strain on the garbage collector and improving performance. Contrary to NaN-tagging, which unboxes all floats, self-tagging adds no overhead to the implementation of other types. Since it does not rely on the specificity of the IEEE754 NaN encoding, self-tagging is highly portable, including to 32-bit architectures.
It is also straightforward to retrofit in implementations that already use object tagging. The only required changes are to reserve one or more tags to self-tagging and implement the encoding and decoding of the target type. The rest of the runtime stays unchanged.
Self-tagging is a good candidate for languages such as JavaScript, that rely heavily on floats, but has potential applications in any polymorphic language.