### v2

# Float Self-Tagging

Abstract.
Dynamic and polymorphic languages must attach information, such as types, to run time objects, and therefore adapt the memory layout of values to include space for this information. It makes it difficult to efficiently implement IEEE754 floating-point numbers as this format does not leave an easily accessible space to store the type information. The two main floating-point number encodings in-use to this day, tagged pointers and NaN-boxing, have drawbacks. Tagged pointers entail a heap-allocation of all float objects, and NaN-boxing puts additional run time costs on type checks and the handling of all other objects.
This paper presents self-tagging, a new approach to object tagging that can attach easily accessible type information to N-bit objects while retaining the ability to use all of their N bits for data. At its core, self-tagging exploits the fact that some bit sequences appear with high probability. Superimposing tags with these frequent sequences allows encoding both N-bit data and type within a single N-bit machine word. The main application of this approach is to represent IEEE754 64-bit and 32-bit floating point numbers on 64-bit and 32-bit machines respectively. We have implemented related variants of self-tagging in one JavaScript compiler and two distinct Scheme compilers to analyze their performance and compare them to tagged pointers and NaN-boxing. Our experiments demonstrate that in practice the approach eliminates heap-allocation of IEEE754 floating-point numbers and improves the execution time of float-intensive benchmarks in Scheme by 2.4$\times$, and in JavaScript by 3.6$\times$, with a negligible performance impact on other benchmarks, which makes it a good alternative to both tagged pointers and NaN-boxing.
Dynamic Languages, Polymorphic Languages, Floating-Point, NaN-Boxing, Compiler Optimization, JavaScript, Scheme
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Object-Oriented Programming, Systems, Languages & Applications; October 12–18, 2025; Singaporeisbn: 978-1-4503-XXXX-X/18/06ccs: Software and its engineering Source code generationccs: Software and its engineering Polymorphism
1. Introduction
In dynamic and other polymorphic languages, efficiently handling floating-point numbers (floats) remains an elusive problem in search of an optimal solution. These languages require attaching type information to values at run time, including numeric types like integers and floats. This requirement conflicts with the IEEE754 standard (iee, 2019) for encoding floats, which uses a full 64-bit word to represent a double-precision float, leaving no space for additional type information. Various representations have been developed to mitigate this issue, each with its own trade offs.
This paper presents a novel technique for encoding floats that avoids most float allocations without adversely affecting the cost of checking types and accessing objects, which are frequent operations of dynamic languages. This technique does not require any program static analysis and could be integrated in most runtime systems of languages that need run time types. It is developed in the context of 64-bit runtime systems but also has applications on 32-bit systems. The rest of this section presents the main encodings in-use as well as their trade offs.
1.1. Tagged Objects
The classical and popular solution to preserve type information is to attach a tag to all objects (Gudeman, 1993; Leijen, 2022; Jones et al., 2012a; Appel, 1998). An N-bit machine word is used to encode an object reference (either a value or a pointer to a heap-allocated value) and a small bit field is reserved in the word to store type information. For the rest of the paper the term object is used interchangeably with object reference. Tagging allows low-cost type checks and object access at the cost of losing a few bits for data. Aligning all heap-allocated values to 64-bit machine words conveniently frees the low bits of pointers to store a 3-bit tag. On many architectures, accessing the fields of the object can be done at no additional cost by using an offset in the memory dereference instruction to take into account the tag for that type of object. Objects can thus be encoded using one of the three following representations.
  • Tagged values store type information with a tag, and the object’s value in the remaining bits. This representation does not require heap allocation.
  • Tagged pointers also store type information with a tag, but values are stored in the heap. The objects are represented by pointers with a tag in their low bits.
  • Generic pointers represent all remaining object types as pointers with a generic tag. Since the generic tag is associated with multiple types, type information is stored in the heap, typically in a header.
Hence, a tag indicates a type, but also a memory representation. A tagged value representation is the most efficient as it needs no heap allocation and no memory read. A tagged pointer still allows efficient type checks, but requires heap allocation. This introduces the overhead of dereferencing the pointer, and adds a strain on the garbage collector. A generic pointer is the least efficient representation as it involves storing type information in the heap, which increases space usage and the cost of type checks and memory management. Consequently, specific tags are a precious resource that must be carefully assigned to the most frequent types observed in programs, such as small integers or floats. Figure 1 shows a typical memory layout of each representation.
Figure 1. The three representations in a tagged object system (here shown on a little-endian machine)
As an example, consider an implementation using 3-bit tags. Due to the frequent use of small integers, a tagged value representation is appropriate for small integers and the choice of the tag 000 allows addition/subtraction directly on the tagged values. Since 3 bits are occupied by the tag, small integers are limited to 61 bits. This is generally considered an acceptable trade off since a two’s complement representation of signed integers can accommodate any number of bits, and 61 bits still offers a wide range of values. When larger integers are required, generic pointers to heap-allocated big integers can be used (Leijen, 2022). Consequently this provides a hybrid representation where the most common integers appear as efficient tagged values and less-frequent big integers are represented with more costly generic pointers.
Unfortunately, when it comes to the representation of floats, none of the aforementioned options are well-suited. Double-precision floats cannot be represented as tagged values since the IEEE754 standard enforces the use of 1 sign bit, 11 exponent bits and 52 mantissa bits (see Figure 2(a)), totaling 64 bits and leaving no space for a tag (iee, 2019). Parting from this standard is impracticable due to the set of instructions offered by x86 (Corporation, 2024) and ARM (Limited, 2009) architectures being specific to 32 or 64-bit floats.
Yet, tagged pointers introduce two major inefficiencies for floats. First, all uses of the value require fetching it from memory. Second, the heap allocation of all floats increases memory usage and the cost of computing float results. This is especially disconcerting considering that allocated floats are often short-lived intermediate results that add a strain on garbage collectors (Serrano and Feeley, 1996; Jones et al., 2012b). This has costly implications for programs performing extensive numerical computations or languages such as JavaScript where floats are the only available primitive number type (International, 2015).
Still, tagged objects are straightforward to implement and are thus found in numerous compilers such as V8 (Sheludko and Solanes, 2020), QuickJS (Bellard and Gordon, 2024), and Hopc (Serrano, 2021) (JavaScript), the Lua interpreter (Roberto Ierusalimschy, 2024) (Lua), CRuby (Shaughnessy, 2013) (Ruby), SBCL (Newman et al., 2024) (Lisp), and Gambit (Marc Feeley, 2025) and Bigloo (Manuel Serrano, 2024) (Scheme). Recently, CPython also moved toward tagged pointers in the process of removing its global interpreter lock (Gross, 2023).
1.2. NaN-boxing
NaN-boxing circumvents the drawbacks of tagged pointer floats by reclaiming unused bits in the encodings of floating-point NaN values to store data (Gudeman, 1993; Leijen, 2022). As per the IEEE754 standard, a NaN value is represented by setting all 11 bits of the exponent to 1 and a non-zero mantissa (zero is used to encode Infinity). The mantissa’s highest bit distinguishes between a quiet and signaling NaN, which determines whether the NaN should signal an exception or fall through operations.
A subset of NaN’s can thus be reserved to encode non-float objects. A convenient subset is that of negative, quiet NaN’s, illustrated in Figure 2(b), which correspond to the interval from 0xfff8_0000_0000_0000 to 0xffff_ffff_ffff_ffff. This partitions floating-point numbers in two intervals. Bit patterns above 0xfff8_0000_0000_0000 are reserved for non-float objects. The bit patterns below 0xfff8_0000_0000_0000 are reserved for floats (including signaling NaN but not quiet NaN) and negative quiet NaN are mapped to the bit pattern 0xfff8_0000_0000_0000.
(a) IEEE754 floating-point representation of $\pi$.
(b) NaN-boxing uses the 51 free bits on negative, quiet NaN values to store data.
Figure 2. Floating-point representations of numbers and NaN.
On current hardware, memory addresses typically fit in 48 bits. The 51 bits uncovered by NaN-boxing are thus sufficient for storing tagged pointers with 3-bit tags, and 32-bit small integers. It offers the advantage of unboxed floats, but impacts the performance of all other, non-float objects due to higher-cost machine instructions necessary to check the type and dereference NaN-boxed pointers.
This reliance on hardware specific details also interferes with other optimizations and portability. Moreover, NaN-boxing on 32-bit architectures would severely cripple memory management because 32-bit floats have NaNs that leave space for only 22-bit pointers, which only allows addressing 4 MiB.
NaN-boxing is used in a few language implementations, including SpiderMonkey (Foundation, 2024) (JavaScript), tinylisp (Engelen, 2024) (Lisp), LuaJIT (Pall, 2024) (Lua), and Zag (Mason, 2022) (Smalltalk).
1.3. Pointer-Biased NaN-boxing
NuN-boxing is a variant of NaN-boxing that alleviates the cost of dereferencing NaN-boxed pointers (Leijen, 2022). It relies on the (currently valid) fact that common hardware never returns negative, quiet NaN’s greater than or equal to 0xfffe_0000_0000_0000.
Figure 3. Encoding of Nun-boxed values, after having applied the 0x0001_0000_0000_0000 bias to floats. The low and high ranges (whose 16 highest bits are 0x0000 or 0xffff) are used for non-float objects that can be represented as tagged values or pointers.
Instead of taking advantage of the unused NaN value range from 0xfffe_0000_0000_0000 to 0xffff_ffff_ffff_ffff, which would result in an encoding similar to standard NaN-boxing, NuN-boxing biases all floats by adding 0x0001_0000_0000_0000. This frees up the lowest and highest ranges for 48-bit tagged objects as shown in Figure 3. It however reintroduces a cost for encoding and decoding floats (adding/subtracting the bias) to operate on their value.
NuN-boxing is used by JavaScriptCore (Inc., 2024) (JavaScript).
1.4. Contribution
The contribution of this paper is a thorough analysis of a new approach to object tagging, named self-tagging, that allows attaching type information to some 64-bit objects while retaining the ability to use all of their 64 bits for storing data. This approach is applied to implement double-precision floats as tagged values instead of tagged pointers. Contrary to NaN-boxing, the newly obtained representation does not impact the performance of encoding and decoding other tagged objects. It is also applicable in the context of single-precision floats on 32-bit architectures, which NaN-boxing does not support. Self-tagging cannot represent all floats but the conducted experiments show that the set of numbers it can represent is large enough to cover all floats actually used by all surveyed JavaScript and Scheme test programs. This paper shows that with float self-tagging there are very few if any heap-allocated floats in practice, thus providing an interesting trade off between the advantages of tagging and NaN-boxing.
1.5. Paper Structure
This paper is structured as follows. Section 2 describes the general idea of self-tagging with a focus on its application to floats. Section 3 discusses a basic form of self-tagging. Section 4 shows how to refine the approach to greatly reduce heap-allocation of floats. Section 5 presents technical details of its implementation. It shows portable implementations that could be used by compilers that generate either C or assembly code. Section 6 shows experimental results both in the context of Scheme with the Bigloo and Gambit compilers, and in the Hopc ahead-of-time JavaScript compiler. Section 7 discusses the adaptation of self-tagging to 32-bit architectures. Finally, section 8 presents related work.
2. Self-Tagging
This section describes the basic form of self-tagging, a tagging technique that exploits the fact that some values naturally contain the appropriate tag corresponding to their type, at the correct location, in their bit arrangement. While this may appear coincidental, it occurs frequently enough to allow the transformation of a significant proportion of heap allocated objects (either represented as tagged pointers or generic pointers) to tagged values. In this paper, unless stated otherwise, double-precision floats and a 64-bit architecture are implied. However, self-tagging pertains to 32-bit architectures as well, which is discussed in Section 7.
Consider the common case of tagged objects with 3-bit tags and, for the sake of the example, assume that the tag 001 is chosen for floats. This setup was previously illustrated in Figure 1 where a tagged pointer refers to the heap-allocated float 3.02. The IEEE754 encoding of 3.02 happens to end with 001, the tag for floats in this example. This purely coincidental fact is bound to happen with $1/8$ of all the bit patterns in the floating-point number space, since eight distinct tags exist on 3 bits.
Self-tagging exploits occurrences where the tag bits of a pointer appear in its value. Such objects can be unboxed, making them tagged values instead of heap allocated values. However, since only $1/8$ of all floats can be unboxed in such a way, a second tag must be reserved for the remaining floats, which still need to be heap allocated with either tagged or generic pointers.
Self-tagging introduces the notion of tags that indicate which values are unboxed. This contrasts with standard object tagging that invites assigning tags with efficient handling of frequent types in mind, regardless of values. With self-tagging, the initial choice of 001 for floats ought to be reevaluated. A better choice could be 000, which captures, among other values, all integer valued floats below $12^{50}+$ 1 (the first integer valued float whose IEEE754 representation does not end in 000). The tag 001 can be kept for tagged pointer floats.
Figure 4. Representation of a (unboxed) float as a tagged value with tag 000. All floats whose IEEE754 representation ends with 000 are unboxed. Other floats are encoded as tagged pointers as usual.
Figure 4 shows the encoding of floats whose IEEE754 representation ends with the better suited tag 000 for unboxed floats. In those cases, the last three bits of the mantissa and the tag are identical. This allows superimposing them to represent both a 64-bit value and a 3-bit tag within a single machine word. Any object tagged with 000 is known to be a tagged value float. Obtaining its value requires no further transformation.
This specific example of self-tagging allows unboxing $1/8$ of floats. Yet, this proportion can be increased by assigning extra tags to unboxed floats. In general, if $T$ tags are available and $f$ tags are reserved for self-tagging of floats, then $f/T$ of all floats can be unboxed. A higher number of reserved tags $f$ can potentially lead to less memory allocation, and thus less strain on the garbage collector. However, the short supply of tags suggests that $f$ should be limited and only worthwhile tags should be selected. Tag selection is discussed in Section 4.
3. A First Experiment
This section presents the first experiment of float self-tagging with the Bigloo Scheme compiler (Manuel Serrano, 2024). Bigloo uses tagged objects by default and can be configured to use NaN-boxing and NuN-boxing (discussed in Section 6.3), making it ideal for comparing these techniques and self-tagging.
3.1. Overview of the Bigloo Compiler
The Bigloo compiler builds a typed abstract syntax tree from the compiled Scheme program. Several analysis and optimization refine the initial optional annotations contained in the source program, the main ones being occurrence-typing (Tobin-Hochstadt and Felleisen, 2010) and storage use analysis (Serrano and Feeley, 1996). During compilation, primitive values have a dual representation: one for polymorphic contexts and one for specific contexts. For instance, when it can establish that a function is always invoked with floats, the compiler assigns the C double type to that function and, when needed, it introduces cast operations from and to the generic representation of floats and double.
The result of this compilation technique is that most temporary and local functions are precisely typed and avoid polymorphic representations, which contributes to removing otherwise necessary heap allocations and casts for floats. The compiler also tracks cases where only floats are stored into vectors, which are then transformed into vectors of C doubles. The experiment conducted in this section evaluates the impact of removing the memory allocations of floats that the compiler is not able to remove statically. For a compiler that does not implement such static analyses, the impact might be even more significant than measured with Bigloo.
3.2. Experimental Setup
Two versions of the Bigloo Scheme compiler are compared. The first is the original Bigloo version, which implements standard tagged objects where floats are heap-allocated using a tagged pointer. The second implementation only differs by its use of self-tagged floats.
In the self-tagging variant, the two tags 000 and 100 are reserved for self-tagged floats, thus increasing the proportion of self-tagged floats to $1/4$. This means that all floats $n$ that are integers such that $|n|\leq 2^{51}$ as well as those values scaled by a power of two, will be unboxed floats. The tag 001 is reserved for the remaining floats, which are represented by tagged pointers.
Both implementations are tested using a subset of the R7RS benchmarks, which is the standard benchmark suite for Scheme (r7r, 2024). The experiment includes all macro-benchmarks (more than 500 lines of code) and benchmarks whose calculations mainly involve floats (tagged as such in the R7RS benchmarks suite). For the purpose of analyzing results, it is useful to distinguish between benchmarks that use few or no floats and float-intensive benchmarks (fibfp=doubly recursive Fibonacci with floats, fft=1024 point Fast Fourier Transform, mbrot=mandelbrot set, nucleic=3D structure determination of a nucleic acid, pnpoly=determine if a 2D point is in a 20 sided polygon, ray=raytrace a scene with 33 spheres, simplex=linear programming using simplex method, sum1=add a file of floats, and sumfp=add the floats 0 to $10^{6}$).
The following results were obtained on a machine with an Intel Core i7-7700K, 48 GB of RAM, and under Debian 10.13 with kernel version SMP Debian 4.19.269-1.
3.3. Memory Allocation
(a) Relative memory allocation
(b) Relative execution time
(c) Relative missed branch prediction
Figure 5. Heap allocations, relative execution time and missed branch predictions comparison between the original Bigloo version and the self-tagging version. The y-axis shows each measure relative to the original version. 1.0 indicates no change. For all metrics, lower is better.
Memory usage is measured by counting the number of heap-allocated bytes across the lifetime of each benchmark. The original and self-tagging Bigloo versions are then compared to assess the efficiency of self-tagging at unboxing floats and its impact on benchmarks that make limited use of floats. Results are shown in Figure 5(a).
On all benchmarks extensively using floats, memory allocation decreases. On mbrot, nucleic, simplex and sum1, allocations decrease by about 25%, which is consistent with the proportion of self-tagged floats ($1/4$). The fft benchmark has only 17% fewer allocations due to predominantly using the float 0.0, which the original Bigloo version preallocates. For fibfp, pnpoly, ray and sumfp there is very little allocation of floats when using self-tagging because those programs mostly use “simple” floats that are integer valued floats $\leq 2^{51}$ scaled by a small power of two, such as 0.75, 2.25, or 42.0.
On all benchmarks that use few floats, memory usage increases by less than 1%. This tiny increase stems from the use of generic pointers for types that were assigned specific tags in the original version.
3.4. Execution Time
Figure 5(b) shows the execution time comparison between both implementations. Execution time refers to elapsed real time measured with perf-stat. Each benchmark is parameterized for a run to last for at least 10 seconds, and is repeated 5 times. All timing results are the average execution time of these runs.
On a geometric average, float-intensive benchmarks are about 25% faster while non-float benchmarks are unaffected. However, the relation between lower memory usage and faster execution is noisy. The execution of mbrot and nucleic is significantly slower despite allocating about $1/4$ less floats with self-tagging.
This unpredictability is explainable, and fixable. Figure 5(c) shows that, despite reducing float allocations in mbrot and nucleic, self-tagging also causes about 4 times more missed branch predictions. This stems from the fact that tags occupy the low bits of the mantissa. In that position, the smallest variations cause a float’s representation to switch between tagged value and tagged pointer in a hard to predict way. This is beyond the capabilites of the branch predictor and execution suffers costly misprediction penalties (Bryant and O’Hallaron, 2016).
5 most 5 most
signif. signif.
expo. expo. Value
bits bits + 1 range fft fibfp mbrot nucleic pnpoly ray simplex sumfp sum1
0.0 94% 13% 0% 0% 16% 13% 11% 0% 0%
00000 00001 5e-324 .. 2.1e-289 - - - - - - - - -
00001 00010 2.1e-289 .. 3.8e-270 - - - - - - - - -
00010 00011 3.8e-270 .. 7e-251 - - - - - - - - -
00011 00100 7e-251 .. 1.3e-231 - - - - - - - - -
00100 00101 1.3e-231 .. 2.4e-212 - - - - - - - - -
00101 00110 2.4e-212 .. 4.4e-193 - - - - - - - - -
00110 00111 4.4e-193 .. 8.1e-174 - - - - - - - - -
00111 01000 8.1e-174 .. 1.5e-154 - - - - - - - - -
01000 01001 1.5e-154 .. 2.8e-135 - - - - - - - - -
01001 01010 2.8e-135 .. 5.1e-116 - - - - - - - - -
01010 01011 5.1e-116 .. 9.4e-97 - - - - - - - - -
01011 01100 9.4e-97 .. 1.7e-77 - - - - - - - - -
01100 01101 1.7e-77 .. 3.2e-58 - - - - - - - - -
01101 01110 3.2e-58 .. 5.9e-39 - - - - - - - - -
01110 01111 5.9e-39 .. 1.1e-19 - - - - - - - - -
01111 10000 1.1e-19 .. 2 6% 33% 96% 70% 60% 18% 62% 0% 0%
10000 10001 2 .. 3.7e19 0% 54% 4% 30% 24% 69% 27% 100% 100%
10001 10010 3.7e19 .. 6.8e38 - - - - - - - - -
10010 10011 6.8e38 .. 1.3e58 - - - - - - - - -
10011 10100 1.3e58 .. 2.3e77 - - - - - - - - -
10100 10101 2.3e77 .. 4.3e96 - - - - - - - - -
10101 10110 4.3e96 .. 7.9e115 - - - - - - - - -
10110 10111 7.9e115 .. 1.5e135 - - - - - - - - -
10111 11000 1.5e135 .. 2.7e154 - - - - - - - - -
11000 11001 2.7e154 .. 4.9e173 - - - - - - - - -
11001 11010 4.9e173 .. 9.1e192 - - - - - - - - -
11010 11011 9.1e192 .. 1.7e212 - - - - - - - - -
11011 11100 1.7e212 .. 3.1e231 - - - - - - - - -
11100 11101 3.1e231 .. 5.7e250 - - - - - - - - -
11101 11110 5.7e250 .. 1.1e270 - - - - - - - - -
11110 11111 1.1e270 .. 1.9e289 - - - - - - - - -
1.9e289 .. 1.8e308 - - - - - - - - -
11111 00000 Infinity/NaN - - - - - - - - -
Figure 6. The positive value ranges captured by each combination of the 5 most significant exponent bits and the proportion of all floats computed by the float-intensive programs in the R7RS benchmark suite. Intervals of interest are highlighted with a background color. Empty entries indicate no float in this range was computed, whereas 0% means that very few floats in this range were generated (less than 0.5%).
4. Improved Predictability and Coverage
This section addresses the issue uncovered by the previous experiment and presents an improved, performant variant of float self-tagging.
The previous experiment shows that while self-tagging decreases the number of heap-allocated floats it does not consistently improve performance due to failed branch predictions. To address this issue, self-tagging is used for the most common floats encountered in practice to maximize the likelihood that a float will be self-tagged, thus improving branch prediction and decreasing heap-allocated floats.
But which floats are most common in programs? To answer this question the float-intensive programs of the R7RS benchmark suite are instrumented to create a profile of all the floats computed by those programs. The results are shown in Figure 6. The table contains 32 rows, one for each combination of the 5 most significant bits of the float’s 11-bit exponent field, which is indicated in the first column. For now ignore the second column. Each of the right columns indicates what proportion of computed floats have a specific combination of most significant exponent bits. The computed floats are clearly clustered around 2, specifically in the range $1.1\times 10^{-19}$. $3.7\times 10^{19}$, and the value 0.0 also occurs frequently for some programs. It is clear that other scientific programs may exhibit profiles that use other ranges of floats. In particular, Infinity and NaN may show up more often in some types of computations. It is apparent that many programs will operate with floats that are predominantly those around 2 on a logarithmic scale, and also 0.0, Infinity, and NaN.
4.1. Self-Tagging with 3 Tags and 4 Tags
Figure 7. A float self-tagging representation where the tag corresponds to the high bits of the exponent. The top bit sequence is a float where the tag is superimposed with the exponent’s three most significant bits. The bottom sequence is the tagged value representation of the float where a 4-bit left rotation is applied to place the tag on low bits.
The float profile of Figure 6 has the nice property that all floats computed have the 3 most significant bits of the exponent field that are either 000 (green rows), 011 (blue rows), or 100 (orange rows). Moreover, among the values captured by 000, only $\pm$0.0 are used.
These 3 exponent bits can be shifted using a bit rotation to align them with the lower 3 tag bits as shown in Figure 7. Thus a float is encoded to a tagged value by a 4-bit rotation to the left (or 60-bit to the right). A tagged value is decoded to a float with the inverse rotation.
Contrarily to the implementation of self-tagging in Section 3, these tags capture intervals of contiguous values. This contiguity within intervals prevents small variations from causing switches from a self-tagged value to a heap-allocated float.
Test programs did not compute the float values $\pm$±Infinity and NaN, but those values would be assigned the tag 111 (gray rows). Assigning this fourth tag to self-tagged floats might be useful to avoid heap-allocated floats in programs frequently computing these values.
Given the expected rarity of the remaining floats there isn’t much of an advantage to represent them using tagged pointers. Instead a representation using a generic pointer with a header type is a better tradeoff as it allows using the tag that would have been used for some other, more frequently used type.
4.2. Self-Tagging with 2 Tags
As is shown in the float profile of Figure 6 the tag 000 is only used to capture the values $\pm$0.0. It is possible to avoid using the tag 000 for self-tagging by preallocating the values $\pm$0.0 using the generic pointer representation. When a float must be encoded as an object it is compared to zero, in which case one of the preallocated zeroes is returned, thus avoiding a new allocation. The same can be done for the values $\pm$±Infinity and NaN, which are expected to be the most frequent values tagged with 111. If this is done then only 2 tags are needed to cover the most frequent cases: 011 and 100.
Unfortunately, this approach may raise the number of failed branch predictions due to the test for the special values $\pm$0.0, $\pm$±Infinity, and NaN. Consequently it may be a more performant tradeoff to avoid testing for some of the special values which are instead handled with a heap-allocation. The best choice will depend on the context, for example the float profile of Figure 6 would suggest not testing for $\pm$±Infinity and NaN.
4.3. Self-Tagging with 1 Tag
Self-tagging can also be achieved with a single tag by a simple transformation of the most significant exponent bits. The 5 most significant bits of the exponent are added to 1 and the middle 3 bits are used as the tag (using a rotation 5 places to the left). This corresponds to the second column of Figure 6. Note that the tag 000 (bold pink rows) covers all the useful ranges, including $\pm$±Infinity and NaN that are not covered by the 2-tag and 3-tag variants.
It is noteworthy that there is nothing special about the tag 000. If for some external reason some other tag is more convenient it can easily be assigned to 1-tag self-tagged floats by adding $1+2\times tag$ rather than 1 to the 5 most significant exponent bits.
Putting all of this together, the self-tagged encoding of a float whose IEEE754 64-bit representation is the integer $n$ can be computed as $(n\oplus((1+2\times tag)<<58))<<_{rot}5$, where $\oplus$ is addition modulo $2^{64}$ and $<<_{rot}$ is the left rotation operator. The operations are reversed to perform the decoding of a self-tagged float to its IEEE754 64-bit representation.
This variant of self-tagging has the advantages of covering all the useful ranges of float values and using a single tag, which leaves more tags available for representing other objects. The only downside when compared to the 4-tag variant is that it covers proportionately fewer ranges of floats, so there is a risk of more heap-allocated floats.
4.4. Self-Tagging with a Different Exponent Bias
A small change in the exponent bias of the 1-tag variant can be used to get another self-tagging variant. If $2\times tag_{1}$ is added to the highest 5 bits and the middle 3 bits are examined, then when they are equal to $tag_{1}$ or $tag_{2}=tag_{1}-1$ they are self-tagged floats. This effectively doubles coverage by including both the light and bold pink rows of the second column of Figure 6. In other words the following (absolute value) ranges will be self-tagged: $0..3.8\times 10^{-270}\cup 5.9\times 10^{-39}..6.8\times 10^{38}\cup 1.1\times 10^{270}..Infinity/NaN$. To put this into perspective, this is a superset of the IEEE754 32-bit floats, so any floating point value that can be represented as a IEEE754 32-bit float can be self-tagged with this variant which uses 2 of the 8 3-bit tags. This is very likely a better use of 2 tags as the 2-tag variant of Section 4.2 does not cover the ranges containing 0, Infinity, and NaN which need special treatment that slows down execution.
When compared to the 1-tag variant, there will be no difference in heap allocations for programs with similar float profiles as Figure 6. However, as will be explained in the next section, it may yield more compact and faster code, in essence because the bias can be added after the rotation, and the need for a bias disappears when $tag_{1}$=000.
5. Implementation
This section discusses the implementation of the float-related operations of the dynamic language. The types f64 and f32 will refer to IEEE754 64-bit and 32-bit doubles and the types i64 and i32 will refer to 64-bit and 32-bit integers. Moreover, the type object will refer to the values of the language (the sum type of floats, integers, booleans, and other objects). To better understand the low-level implications of the implementation and give a sense of the execution time performance, assembly code for the Intel x86 architecture is shown to illustrate the details of the implementation on that popular architecture. Although this section focusses on a 64-bit implementation where the type object is an i64 and floats are f64, it is fairly straightforward to implement the operations similarly on a 32-bit implementation where the type object is an i32 and floats are f32.
Testing the low tag bits of an object value is a basic need for float-related operations to convert f64 to and from object and for dynamic type checks. In particular it is possible to determine if a f64 can be self-tagged by going through the encoding process and checking that the resulting tag is one of the tags (or the tag) appropriate for the chosen self-tagging variant.
5.1. Single Tag Testing
Checking that the N low bits are equal to a specific tag can be achieved by masking those bits followed by a comparison. Assuming the object value is in the x86 64-bit register rax, then the following sequence will branch to label tag_matches when the 3 low bits are equal to tag [^1] Intel assembler syntax is used throughout.:
and al, 7
cmp al, $tag$
jz tag_matches
Note that by using al only the lower 8 bits of rax are accessed, leading to smaller constants and instruction encodings. However, this sequence modifies rax so if the value is needed after the tag test, as often will be the case, then an additional instruction and register are required to keep a copy. In the special case of tag=0, the two first instructions can be replaced with “test al, 7”, which is a bitwise-and that tests if all 3 lowest bits of rax are zero without modifying rax. In the general case when $\neq$0 the test instruction can be used by combining it with a lea (Load Effective Address) instruction that copies to a temporary register the source register plus a constant that cancels tag (modulo 8):
lea ebx, [eax+(8-$tag$)]
test bl, 7
jz tag_matches
5.2. Multiple Tag Testing
When it is necessary to check if an object’s tag is one of a set of tags it is often possible to do this more efficiently than a sequence of single tag checks. This is particularly the case when there is leeway in the assignments of tags as is often the case.
[…]
lea ebx, [eax+(8-3)] # 011 -> 000 and 100 -> 001
test bl, 6
jz tag_matches_011_or_100
This technique can be extended when testing the 4 tags 000, 011, 100, and 111 of the 4-tag self-tagging approach:
lea ebx, [eax+(8-7)] # 000 -> 001, 011 -> 100, 100 -> 101, and 111 -> 000
test bl, 2
jz tag_matches_000_or_011_or_100_or_111
[…]
mov bx, 0x1919 # set of tags
bt bx, ax # test bit at index ax of bx
jc tag_matches_000_or_011_or_100
The bit index in ax is obtained mod 16, so the byte 0b00011001=0x19 (all 0’s except at bit indices 0, 3, and 4) is repeated twice in register bx. Using the bt instruction has the advantage of neither modifying the rax register nor the register holding the set of tags, so the cost of the mov instruction can be amortized over multiple checks of that set of tags (a possibly near zero cost if that register is globally reserved).
On architectures without a bit test instruction, a dynamic count shift of the tag set register could be used to achieve the bit indexing. If this is done by modifying the tag set register the initialization can’t be shared by multiple tag tests. On most 3-address RISC architectures the shift can be done non-destructively. On ARM A64, here is a short instruction sequence that tests the 3 low bits of 64-bit register x1 for a tag of 0, 3, or 4:
lslv w3, w2, w1
cmp w3, 0
bmi matching_tag
It assumes that the 32-bit register w2 has been preloaded with 0x98989898, the bit reversed tag set that aligns the bit for tag 0 with the sign bit.
The cost of the check does not depend on the number of tags tested. This can be advantageous if a tagged pointer is used for heap-allocated floats, say with tag h. In that case, a check for a float (either self-tagged or heap-allocated) can be done by adding h to the tag set. Once it is known that an object is a float, it is easy to check for the single tag h to discriminate between the self-tagged and heap-allocated representations.
[…]
To convert a f64 to an object, it must first be converted to an i64, and then a constant $bias$ must be added modulo $2^{64}$ followed by a bit-rotation by $r$ places left. For the 1-tag variant, $bias=(1+2\times tag)\times 2^{58}$ and $r=5$. For the 2-tag, 3-tag, and 4-tag variants, $bias=0$ (no bias) and $r=4$. After this there is a check to verify if the resulting object’s tag is in the set of self-tagged tags. If this is the case the conversion is done, otherwise an out of line routine can be called to heap-allocate a float and return an appropriately tagged pointer. As an example, if the f64 value is in the 64-bit float register xmm0, then the following x86 code will set register rax to the corresponding object when the 3-tag self-tagging variant is used:
mov rax, xmm0 # rax $\leftarrow$ xmm0
rol rax, 4 # rotate left 4 places
bt bx, ax # assumes bx initialized to 0x1919 elsewhere
jnc heap_alloc_float # tag is not one of 000, 011, or 100?
done:
Assuming the initialization of the tag mask register (bx) is amortized over multiple float-related operations, the cost for the hot path is low: 3 register-to-register operations and an easily predictable conditional jump.
The following x86 code is appropriate for the 1-tag self-tagging variant using 000 as the tag:
mov rax, xmm0 # rax $\leftarrow$ xmm0
add rax, rbx # assumes rbx initialized to 1<<58 elsewhere
rol rax, 5 # rotate left 5 places
test al, 7
jnz heap_alloc_float # tag is not 000?
done:
Here also, the cost for the hot path is just a bit more due to the bias: 4 register-to-register operations and an easily predictable conditional jump.
For the variant explained in Section 4.4 which uses 2 tags with no special handling of zeroes, it is more efficient to do the rotation before adding the bias because then the bias is the small constant $tag_{1}$, and in the special case of $tag_{1}$=000 the bias vanishes. For example, here is the code when $tag_{1}$=011:
mov rax, xmm0 # rax $\leftarrow$ xmm0
rol rax, 5 # rotate left 5 places
add al, 3 # sufficient to modify the lower 8 bits of rax
bt bx, ax # assumes bx initialized to 0x0c0c elsewhere
jnc heap_alloc_float # tag is not one of 010 or 011?
done:
The add instruction uses a small constant as a bias, contrary to the code for the 1-tag version (large constants typically need to be setup in a register, which is additional work even if it is amortized). Moreover the 1-tag version cannot eliminate the bias, but here the add instruction can be removed when $tag_{1}$=000 which makes it slighly faster than the 1-tag variant and as efficient as the 3-tag variant: 3 register-to-register operations and an easily predictable conditional jump.
[…]
To convert an object (that is known to be a float) to a f64, the tag must be tested to see if it is a self-tagged float. If it is, the boxing operations are done in reverse order and inverted, replacing the rol by a ror (rotate-right), and if there is a bias, replacing the add by a sub (subtract). If it is not a self-tagged float, a memory read gets the f64 result out of the heap allocated float. As an example using the 3-tag variant, if the object value is in the 64-bit register rax, then the following x86 code will set register xmm0 to the corresponding f64:
bt bx, ax # assumes bx initialized to 0x1919 elsewhere
jc self_tagged
mov xmm0,[rax+offset] # xmm0 $\leftarrow$ value field of float
jmp done
self_tagged:
ror rax, 4 # rotate right 4 places
mov xmm0, rax # xmm0 $\leftarrow$ rax
done:
The number and type of instructions on the hot path is the same as for the boxing operation.
5.5. C Implementation
This section explains how the above operations can be implemented in C, as this is a common implementation language, in particular it is used by the Bigloo and Gambit Scheme to C compiler used in the experiments of Sections 3 and 6.
Testing for a specific tag or a pair of tags is fairly straightforward to express in C and follows the same principles as the assembly code, so details are skipped. Testing for a set of tags is more challenging in portable C code. It can be implemented with the following pure C function, which is easily inlined by a C compiler (here testing for the tags 000, 011, and 100):
#define TAG_SET ((1<<0)|(1<<3)|(1<<4)) /* 0x19 for tags 000, 011, and 100 */
inline bool has_tag_0_or_3_or_4(int64_t n) {
return (((uint32_t)1 << (n & 31)) & (~(uint32_t)0/0xff * TAG_SET)) != 0;
The expression n & 31 computes a shift count mod 32 and the bitwise-and therefore tests the bit at index n mod 32 of its second operand. The code uses the compile-time constant ~(uint32_t)0/0xff * TAG_SET that repeats the 8 bit mask TAG_SET 4 times to fill a 32 bit word, giving 0x19191919. This achieves the equivalent of testing the bit at index n mod 8 of TAG_SET.
Using this pure C definition with GCC version 13.2.0 and clang version 18.1.3 on x86-64 results in both cases in the non-destructive approach based on the bit test instruction. The expression n & 31 is optimized out by those compilers because the shift instruction on x86 implicitly does this operation. On older versions of those compilers that would implement this with a destructive shift instruction, the bit test approach can be achieved using the following definition that uses an asm statement:
inline bool has_tag_0_or_3_or_4(int64_t n) {
bool carry;
__asm__("bt %%ax, %2;" /* get one bit of mask into carry (ax is index) */
: "=@ccc"(carry)
: "a"((uint16_t)n),
"r"(~(uint16_t)0/0xff * TAG_SET)); /* 0x1919 */
return carry;
Boxing and unboxing operations need to convert between a float value and its bit representation. This can be achieved portably with a union type whose fields are of type f64 and i64. Moreover, boxing and unboxing need to rotate the bit representation of the float. Although C does not provide an operator for bit rotation, both GCC and clang recognize an equivalent pair of shifts and generate a single machine rotate instruction.
The implementation of boxing in C for the 3-tag variant is:
union di { double d; int64_t i; };
#define ROTL(n,s) ((int64_t)(((uint64_t)n << s) | ((uint64_t)n >> (64 - s))))
inline int64_t f64_to_object(double f) {
int64_t result = ROTL(((union di)f).i, 4);
if (has_tag_0_or_3_or_4(result)) return result;
return heap_allocate_float(f);
The implementation of unboxing for the 3-tag variant is:
inline double object_to_f64(int64_t o) {
int64_t result = ROTL(o, 60);
if (has_tag_0_or_3_or_4(result)) return ((union di)result).d;
return value_of_heap_allocated_float(o);
6. Experiments of Self-Tagging using the Exponent Bits
The experimental setup from Section 3 is replicated to test float self-tagging using high bits of the exponent. Three variants are compared to the original version of Bigloo that implements standard tagged objects.
  • 3-tag self-tagging uses the tags 000, 011 and 100, thus self-tagging floats in the green, blue, and orange ranges from Figure 6. The remaining floats are represented by generic pointers with a header type, and thus are heap-allocated.
  • 2-tag self-tagging tests the impact of expending the tag 000 for self-tagging. It only uses the tags 011 and 100 for float self-tagging and reclaims the tag 001 to represent all remaining floats as tagged pointers. The floats $\pm$0.0 are preallocated and the float boxing operation contains an explicit zero check to return the preallocated zeros.
  • 1-tag self-tagging tests the 1-tag variant where the tag 001 is used for self-tagged floats.
Figure 8 shows the memory allocations of these variants. All three variants allocate fewer floats on all float-intensive benchmarks. In fact, all variants have been verified to allocate no float at all (see Figure 6) with the only exceptions being the preallocated $\pm$0.0 by the 2-tag variant. The remaining heap usage is for non-float objects. Similarly to the previous experiment (Figure 5(a)), memory usage of other benchmarks remains mostly unaffected.
Figure 8. Heap allocation comparison between the three variants of self-tagging on exponent bits. The 3-tag variant uses the tags 000, 011 and 100. The 2-tag variant uses the tags 011 and 100, but preallocates 0.0 and -0.0. The 1-tag variant uses 3 bits of the biased exponent (see Figure 6). The y-axis shows heap allocations relative to the original Bigloo version, which uses tagged pointers for all floats. 1.0 indicates no change, while 0.0 indicates that all allocations were removed.
Figure 9 shows faster execution times for all three variants of self-tagging. All float-intensive benchmarks execute faster than the original Bigloo implementation, with the exception of fft on the 2-tag variant. In particular, the mbrot and nucleic benchmarks are both faster, confirming that self-tagging on exponent bits eliminates the branch misprediction penalty incurred by self-tagging on mantissa low bits (see Section 3). Despite small individual variations, other benchmarks remain generally unaffected.
On the float-intensive benchmarks, execution time with the 3-tag variant, which makes use of the tag 000, is on average 0.41$\times$ that of the original Bigloo version, slightly more than two times faster. This is significantly lower than the 2-tag variant (0.48$\times$) which avoids the tag 000. In particular, the benchmark fft, which predominantly uses the values $\pm$0.0, is slower in the 2-tag variant despite zeros being preallocated. The 1-tag variant (0.42$\times$) is similar to 3-tag on float benchmarks, and both variants display no performance change on non-float benchmarks (1.01$\times$ for 1-tag, and 1.03$\times$ for 3-tag). Due to the 1-tag variant not using 000, and leaving more tags available, it is a more versatile candidate and is thus used as the default variant for the rest of this paper.
Figure 9. Execution time comparison between the three variants of self-tagging on exponent bits. The 3-tag variant uses the tags 000, 011 and 100. The 2-tag variant uses the tags 011 and 100, but preallocates 0.0 and -0.0. The 1-tag variant uses 3 bits of the biased exponent (see Figure 6). The y-axis shows execution time relative to the original version. 1.0 indicates no change, lower means faster execution than the original Bigloo version, which uses tagged pointers.
6.1. Effect on Garbage Collection
Execution time (s)
fft Refer to caption
fibfp Refer to caption
mbrot Refer to caption
nucleic Refer to caption
pnpoly* Refer to caption
ray Refer to caption
simplex Refer to caption
sum1 Refer to caption
sumfp Refer to caption
Preallocated Heap Data Size
Figure 10. Execution time of float-intensive R7RS benchmarks with extra live data allocated on heap before execution. Benchmarks are executed in unsafe mode (without typechecks) to specifically gauge the impact on garbage collection. The x-axis shows the size of preallocate data in bytes on a logarithmic scale. An extra point is added at $x=0$ for execution time without preallocated data. The y-axis represents execution time in seconds. Dashed lines represent systems that do not implement self-tagging (original Bigloo, original Gambit, and Chez Scheme). Solid lines are Bigloo and Gambit with self-tagging (1-tag).
*pnpoly’s execution time with Bigloo (orig) is about 230 seconds, which is cropped out for readability.
Float self-tagging improves performance by allocating fewer floats on the heap, thus relieving pressure on the garbage collector. Hence, the impact of self-tagging is more pronounced for programs that spend more time doing memory management tasks. This section presents an experiment to measure execution time on float-intensive R7RS benchmarks when the heap contains live data, thus putting more pressure on the garbage collector.
The R7RS benchmarks suite is modified to allocate a vector before the execution of each benchmark. The size of the allocated vector is given as a parameter of the benchmark. Each benchmark is then executed normally with vectors of size ranging from $10^{4}$ to $10^{8}$, which correspond to about 80 kB and 800 MB of memory respectively. An execution is also done with no preallocated vector. To specifically measure the impact of self-tagging on garbage collection, benchmarks are executed in unsafe mode, which allows unsafe optimizations such as the removal of safety typechecks.
For this experiment the float self-tagging approach has been implemented in the Gambit 4.9.6 (Marc Feeley, 2025) optimizing Scheme compiler. The benchmarks are executed with Bigloo and Gambit, with and without self-tagging (1-tag), and also the Chez Scheme 10.1 (R. Kent Dybvig, 2025) optimizing Scheme compiler without self-tagging. These compilers are chosen due to their status as state-of-the-art Scheme implementations and the fact they employ distinct garbage collection techniques. Chez Scheme uses a generational garbage collector. Gambit uses a stop-and-copy garbage collector with bump allocation. Bigloo uses the Boehm-Demers-Weiser Garbage Collector (Boehm and Weiser, 1988).
Figure 10 compares execution times of these Scheme implementations. With a nearly empty heap, and thus little strain on the garbage collector, the bump allocator of Gambit without self-tagging generally outperforms all other implementations since it does not incur the cost, albeit small, of encoding and decoding self-tagged floats. However, once the heap is preloaded with 1 to 10 MB (which is roughly the size of the data cache) the strain on all garbage collectors starts to outweigh the cost of encoding and decoding self-tagged floats, thus Bigloo and Gambit with self-tagging hereafter outperform all implementations without self-tagging.
The threshold of 1 to 10 MB of preallocated data beyond which self-tagging increasingly outperforms other implementations suggests that experiments in other sections are in fact worst-case scenarios. On real life programs, where it is reasonable to expect some persistent data in the heap, self-tagging can be expected to offer a more substantial improvement.
6.2. Comparison to NaN-Boxing and NuN-Boxing
Bigloo also contains implementations of NaN-boxing and NuN-boxing which can be enabled through a configuration option.
Figure 11 compares Bigloo’s original representation (tagged pointers) to NaN-boxing, NuN-boxing, and 1-tag float self-tagging. On float-intensive benchmarks, the average execution time with self-tagging (0.42$\times$ of original version) is higher than with NaN-boxing (0.30$\times$) and NuN-boxing (0.34$\times$). On non-float benchmarks the average relative execution time of self-tagging (1.01$\times$), NaN-boxing (1.03$\times$), NuN-boxing (1.03$\times$) all remain similar to the original version. However, on individual benchmarks the performance of self-tagging remains generally unchanged while significant variations are observed with NaN/NuN-boxing.
Figure 11. Execution time comparison between Bigloo with 1-tag self-tagging, Bigloo with NuN-boxing, and Bigloo with NaN-boxing. The y-axis shows execution time relative to the original Bigloo version, which uses tagged pointers. 1.0 indicates no change, lower means faster execution than the original version.
6.3. Application to JavaScript
This section experiments with the application of float self-tagging in JavaScript, which is a prime candidate for self-tagging because it has no integer type. Its specification defines numbers as double-precision floats (International, 2015). To isolate the impact of the representation of floats, four versions of the Hopc JavaScript compiler are tested, where each version differs only by its representation of floats, which is either tagged pointers, NaN-boxing, NuN-boxing, or self-tagging.
Hopc is an ahead-of-time JavaScript compiler that uses Scheme as its compilation backend and relies on the underlying Scheme compiler for optimizations (Serrano, 2021). This permits reusing Bigloo’s runtime to test float self-tagging in the context of JavaScript.
Figure 12 presents the results of porting Bigloo’s NaN-boxing, NuN-boxing, and self-tagging implementations to JavaScript. Performance is measured using the Hop benchmark suite, which gathers benchmarks from the popular JetStream, Octane and Sunspider JavaScript benchmark suites (Serrano, 2024). Once again, it is useful to categorize benchmarks between float-intensive benchmarks (almabench, n-body, navier-stokes, and raytrace), and benchmarks that use few or no floats.
On the float-intensive benchmarks, the relative execution times of Hopc NaN-boxing (0.26$\times$), NuN-boxing (0.26$\times$), and self-tagging (0.28$\times$) implementations are comparable, roughly 4 times faster than tagged pointers. Execution time with NaN-boxing and NuN-boxing is faster on almabench and n-body, but slower on raytrace and navier-stokes. However, on non-float benchmarks the self-tagging implementation (1.03$\times$) generally fares better due to its lower overhead than NaN-boxing (1.12$\times$) and NuN-boxing (1.16$\times$) when handling non-float objects, by up to a factor of 2.07$\times$ for crypto-sha1.
Figure 12. Execution time comparison between Hopc with 1-tag self-tagging, Hopc with NuN-boxing, and Hopc with NaN-boxing. The y-axis shows execution time relative to the original Hopc version, which uses tagged pointers. 1.0 indicates no change, lower means faster execution than the original version.
4 most 4 most
signif. signif.
expo. expo. Value
bits bits + 1 range
0.0
$0$ 0001 1.4e-45 .. 3.9e-34
0001 0010 3.9e-34 .. 2.5e-29
0010 0011 2.5e-29 .. 1.7e-24
0011 0100 1.7e-24 .. 1.1e-19
0100 0101 1.1e-19 .. 7.1e-15
0101 0110 7.1e-15 .. 4.7e-10
0110 0111 4.7e-10 .. 3.1e-5
0111 1000 3.1e-5 .. 2
1000 1001 2 .. 1.3e5
1001 1010 1.3e5 .. 8.6e9
1010 1011 8.6e9 .. 5.6e14
1011 1100 5.6e14 .. 3.7e19
1100 1101 3.7e19 .. 2.4e24
1101 1110 2.4e24 .. 1.6e29
1110 $1$ 1.6e29 .. 1.0e34
1.0e34 .. 3.4e38
$1$ $0$ Infinity/NaN
Figure 13. The positive value ranges captured by each combination of the 4 most significant exponent bits of the IEEE754 32-bit representation. The rows coloured in bold pink are covered by the 1-tag variant (tag 00) and the rows coloured in light pink are covered by the 2-tag variant (tags 00 and 11).
7. Self-Tagging on 32-bit Architectures
Although previous sections only considered double-precision floats on 64-bit architectures, self-tagging of single-precision floats is a natural adaptation for 32-bit architectures. Indeed, the core idea, which is to superimpose a tag to a sequence of bits likely to appear in practice, can be used for any machine word size. A popular object representation on 32-bit architectures is to use 2-bit tags (4 possible tags) and heap-allocated objects aligned to 32-bit words.
Figure 13 shows the ranges covered by each combination of the top 4 bits of the exponent, which is 8 bits wide in the IEEE754 32-bit representation.
The 1-tag approach described previously can be adapted to the narrower exponent of the IEEE754 32-bit representation field: the self-tagged encoding of a float whose IEEE754 32-bit representation is the integer $n$ can be computed as $(n\oplus((1+2\times tag)<<27))<<_{rot}4$, where $\oplus$ is addition modulo $2^{32}$ and $<<_{rot}$ is the left rotation operator. If the two lowest bits are equal to $tag$ then the float is self-tagged. This will cover all float values in the ranges: $0.0..3.9\times 10^{-34}\cup 3.1\times 10^{-5}..1.3\times 10^{5}\cup 1.0\times 10^{34}..Infinity/Nan$ (bold pink rows of Figure 13 using $tag$=00).
The coverage can be doubled by using an adaptation of the 2-tag variant described in Section 4.4: $0.0..2.5\times 10^{-29}\cup 4.7\times 10^{-10}..8.6\times 10^{9}\cup 1.6\times 10^{29}..Infinity/Nan$ (light pink rows of Figure 13 using $tag_{1}$=00 and $tag_{2}$=11). With this variant, it would be reasonable to reserve one tag for small signed integers in the range $-2^{29}..2^{29}-1$ (-536870912 .. 536870911), and the remaining tag for heap-allocated objects, including the floats that can’t be self-tagged.
8. Related Work
Handling floats is a long-standing issue in dynamic language implementation. In the last decades, little progress has been made improving the encoding of floats, with most implementations either using heap-allocated floats or suffering from the overhead of NaN-boxing on pointers. Thus, more general strategies that use data flow analysis (Henglein, 1992; Serrano and Feeley, 1996; Petersen and Glew, 2012; Lindahl and Sagonas, 2003) and partial evaluation (Würthinger et al., 2017) have been developed to find locations where the type of a value is known across its lifecycle. This allows the generation of specialized code that safely handles fully unboxed, untagged values. Such strategies are not specific to unboxing floats. Rather, they tackle the more general problem of inferring types in dynamic languages, which allows dropping type information in either tagged pointers or NaN-boxed pointers. While more limited in scope, self-tagging solves the problem of heap-allocated floats with a new encoding that involves no program analysis.
Self-tagging is straightforward to add to compilers that already use object tagging. Therefore, implementations that represent floats as tagged pointers could benefit from it with minimal implementation effort. Such popular implementations include CPython (Gross, 2023) and Google’s V8 (Sheludko and Solanes, 2020). Compilers that use NaN/NuN-boxing, such as Mozilla’s SpiderMonkey (Foundation, 2024) and Apples’s JavaScriptCore (Inc., 2024) can also benefit from the lower impact that self-tagging has on the performance of non-float types.
Due to the shortcomings of tagged pointers and NaN-boxing, additional strategies are used to avoid boxing. A straightforward approach is to provide homogeneous data structures such as TypedArray in JavaScript (International, 2015) and numpy’s float64 arrays in Python (Harris et al., 2020). Type homogeneity then allows unboxing values within the collection. More generally, compilers can detect data homogeneity at run time and use context-dependent storage strategies to store data in a way that prevents pointer chasing (Bolz et al., 2013). Since self-tagging unboxes most floats in practice, it effectively prevents such pointer chasing in the case of floats without static or run time analysis.
Even for general-purpose languages not specifically designed for numerical computation, ensuring reasonably fast floating-point operations is a major concern. All modern production-quality implementations surveyed for this paper devote significant effort to optimizing them. Unfortunately the academic literature is scarce when it comes to methods and techniques used by these efficient implementations.
In many cases, the source code itself has to be studied to understand the implementation and compare it to self-tagging. The NuN-boxing implementation used in experiments in this paper is based on that of JavaScriptCore [^2] https://github.com/WebKit/WebKit/commit/74c9b9bbc2c7527144ac366c3f62f84aaa1a8a4e. It uses the pattern 0x0000 in the highest 16-bits to encode pointers and 0xffff to encode integers. The other patterns are used for floats. The actual types of heap-values is either encoded in the three lowest bits of the pointer itself or in a header word. For instance, as strings are frequent data type, the pattern 110 is used to encode these values and checking that a value is a string requires to check the lowest three bits and the 16 highest bits. NuN-boxing has the benefits over self-tagging that it encodes all floats uniformly and does not reserve tags for them. Its drawbacks are that it negatively impacts the type checking and accesses of all objects, it can represent integers with at most 48 bits (usually just 32 bits for practical reasons), and it can’t be used on 32-bit architectures.
The technique closest to self-tagging discovered among surveyed implementations are those of CRuby [^3] https://github.com/ruby/ruby/commit/b3b5e626ad69bf22be3228f847f94e1b68f40888 and OpenSmallTalk [^4] https://clementbera.wordpress.com/2018/11/09/64-bits-immediate-floats. Both implementations use a form of self-tagging with a rotation of the mantissa close to the encoding of Section 4.3. However, bit patterns captured by both implementations exclude ±$\pm 0.0$, which is treated as a special case. This imposes complex sequences of instructions for encoding and decoding that can hardly be inlined in the generated code and risks increasing the number of branch misses as demonstrated in Section 3.4.
To the best of the authors’ knowledge, this is the first peer-reviewed paper exposing the principles of self-tagging and evaluating its benefits, drawbacks, and possible variants.
In this paper, self-tagging was discussed primarily in the context of dynamic languages. Yet, polymorphic languages such as OCaml and Haskell also attach information to run time values and spend considerable effort to find efficient encodings for abstract data types (Baudon et al., 2023; Kennedy and Vytiniotis, 2012). Self-tagging could be applied in these languages, either to floats or to the encoding of abstract data types.
9. Conclusion
Avoiding heap allocation of floating point numbers has always been a major concern for dynamic and polymorphic languages. This paper presents and evaluates the popular techniques deployed to that end and it presents a new approach, coined self-tagging, that allows attaching type information to 64-bit values at low cost. The core idea is to superimpose tags to bits that appear with high probability in practice. This approach is especially effective to represent IEEE754 floating-point numbers, whose encoding requires all bits of a 64-bit machine word.
Float self-tagging avoids heap-allocation of frequently occurring float values, reducing the strain on the garbage collector and improving performance. Contrary to NaN-boxing and NuN-boxing, which are designed to optimize the handling of floats, self-tagging offers a different trade off that adds no overhead to the handling of other types. Since it does not rely on the specificity of the IEEE754 NaN encoding, self-tagging is highly portable, including to 32-bit architectures.
It is also straightforward to retrofit in implementations that already use object tagging. The only required changes are to reserve one or more tags to float self-tagging and implement the encoding and decoding of the target type. The rest of the runtime stays unchanged.
As shown by the experimental evidence presented in the paper, float self-tagging is a good candidate for languages such as JavaScript, whose specification relies heavily on floats, but has potential applications in any polymorphic language.
Given the purely bit-shuffling nature of float self-tagging boxing and unboxing, it may also be a good candidate for an essentially zero-cost implementation at the hardware level when floats are moved between float and integer registers.